Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates: Machine-hours required to support estimated production Fixed manufacturing overhead cost Variable manufacturing overhead cost per machine-hour 158,000 652,000 $ 4.40 Required: 1. Compute the plantwide predetermined overhead rate. 2. During the year, Job 400 was started and completed. The following information was available with respect to this job Direct materials Direct labor cost Machine-hours used 360 210 37 Compute the total manufacturing cost assigned to Job 400. 3. If Job 400 includes 50 units, what is the unit product cost for this job? 4. If Moody uses a markup percentage of 110% of its total manufacturing cost, then what selling price per unit would it have established for Job 400? Complete this question by entering your answers in the tabs below. Required Required 2Required 3 Required 4 If Moody uses a markup percentage of 110% of its total manufacturing cost, then what selling price per unit would it have established for Job 400? (Round your intermediate calculations and final answer to the nearest whole dollar.) Selling price per unit
Answer #1

Answers

  • All working forms part of the answer
  • Requirement 1

Working

Molding

A [given]

Total Fixed manufacturing Overhead cost

$652,000

Allocation base

machine hours

B [given]

Amount of allocation base

                 158,000

C = A/B

Fixed Overhead rate

$                     4.13

D [given]

Variable overhead rate

$                     4.40

E = C+D

Predetermined Overhead rates

$                     8.53

  • Requirement 2

Molding

Direct Material Cost

$360

Direct Labor Cost

$210

Manufacturing Overhead cost applied

$316

Total manufacturing cost assigned

$886

  • Requirement 3

Unit product cost = $ 886/50 units = $ 17.72 = Answer

  • Requirement 4

Selling Price per unit = 17.72 + (17.72 x 110%) = $ 38 = Answer